OpenStudy (anonymous):

What is the solution of the proportion? 14/12= d/48 A. 56 B. 672 C. 168 D. 576

OpenStudy (tkhunny):

Multiply the left fraction by \(\dfrac{4}{4}\) and see if it makes any more sense.

OpenStudy (anonymous):

you just have to multiply both side by 48 14//12 = d/48 => (14/12) *48 = (d/48)* 48 => 14 * 4 = d => d = 56 so ans is A. 56